To move a refrigerator of mass 140 kg into a house, a mover puts it on a dolly and covers the steps leading into the house with a wooden plank acting as a ramp. The plank is 23 m long and rises 4 m. The mover pulls the dolly with constant velocity and with a steady force 620 N up the ramp.

1)How much work does he perform? The acceleration of gravity is 9.8 m/s 2 . Answer in units of J.

2)What is the minimal work required to lift the refrigerator into the house? Answer in units of J.
Answer #1

work done by mover in the direction of displacement,

W = Fnet*∆X = F - m*g*sin¥ = (620-140*9.81*sin¥)*23

= {620-140*9.81*(4/23)}*23= 8766.4 Joule

work required to loft refrigerator,

W = m*g*∆Y = 140*9.81*4 = 5493.6 Joule